# Towards Temporally Consistent Referring Video Object Segmentation

### v1

@@ Abstract
Referring Video Object Segmentation (R-VOS) methods face challenges in maintaining consistent object segmentation due to temporal context variability and the presence of other visually similar objects. We propose an end-to-end R-VOS paradigm that explicitly models temporal instance consistency alongside the referring segmentation. Specifically, we introduce a novel hybrid memory that facilitates inter-frame collaboration for robust spatio-temporal matching and propagation. Features of frames with automatically generated high-quality reference masks are propagated to segment the remaining frames based on multi-granularity association to achieve temporally consistent R-VOS. Furthermore, we propose a new Mask Consistency Score (MCS) metric to evaluate the temporal consistency of video segmentation. Extensive experiments demonstrate that our approach enhances temporal consistency by a significant margin, leading to top-ranked performance on popular R-VOS benchmarks, i.e., Ref-YouTube-VOS (67.1%) and Ref-DAVIS17 (65.6%).
@@ I Introduction
Early R-VOS methods [13, 14] primarily employ convolution networks [15] for visual encoding and language grounding. However, their inherent limitations in capturing long-range dependencies and handling free-form features result in suboptimal performance. With the advancement in attention mechanism [16, 17, 18, 19, 20], various attention-based R-VOS approaches have emerged, leading to significant improvements. Techniques such as YOFO [21] and SgMg [22] employ cross-attention to model multimodal representations, enabling flexible and long-range vision-language correspondence. Furthermore, ReferFormer [23] and MTTR [24] introduce end-to-end transformers with object queries to dynamically separate target objects from semantic features.
These transformer-based techniques employ a set of language-guided object queries to attend to multimodal features of different frames, eventually predicting conditional kernels [25] for each frame. The frame-specific conditional kernels are then used to generate object masks by convolving the corresponding multimodal features. Despite demonstrating promising performance, these approaches struggle to maintain consistent segmentation in complex and challenging scenarios. As illustrated in the first row of Fig. 1(b), the target sheep is accurately segmented in the initial frame, but the presence of other similar-looking sheep and changes in the local and global context, e.g., the disappearance of the dog in the final frames, lead to the failure of the model to consistently segment the correct target.
To mitigate this inconsistency issue, some approaches utilize well-trained Semi-VOS models [26, 27] for post-processing, e.g., ReferFormer [23] and PMINet [28] use CFBI [26] as a post-processing step to refine the masks. Such a naïve strategy requires a complex pipeline with additional backbones [15, 29] and training data [30, 31, 32, 33, 34, 35], making R-VOS cumbersome and impractical. Moreover, this strategy does not yield satisfactory improvements in performance as we show in Table IV. An ideal solution would be an end-to-end paradigm that seamlessly integrates temporal consistency into R-VOS.
@@ II-A Semi-supervised Video Object Segmentation
Semi-VOS aims to segment target objects throughout the entire video, given the ground-truth mask (of the first frame) [38, 39, 40, 41]. Existing techniques mainly fall into three paradigms: finetuning-based, tracking-based, and matching-based. Finetuning-based methods dynamically update networks during test time using the first [42, 43], selected [44], or synthetic [45] frames from each video to make models object-specific. Although these methods show promising results, the slow test-time adaptation makes them inefficient. Tracking-based approaches [26, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55] avoid online learning and perform frame-to-frame propagation using masks of neighboring and annotated frames. However, they often lack sufficient long-term context, leading to the loss of targets after occlusions.
To address the context limitations, matching-based methods [56, 57, 58, 59, 60, 61, 62, 63, 12] construct memory banks with features from past frames and use attention mechanisms for temporal propagation. STM [64] and STCN [27] build memory banks to store mask and visual features at fixed intervals. SST [65] uses transformer blocks for affinity computation. AOT [11] and its subsequent works [66] establish hierarchical propagation to process multiple objects in parallel. XMem [10] introduces complementary memories inspired by the Atkinson-Shiffrin model [67] to efficiently handle long videos. In this work, we propose a hybrid memory that computes multi-granularity affinity for robust spatio-temporal propagation, even with imperfect reference masks.
@@ II-B Referring Video Object Segmentation
R-VOS methods conduct vision-and-language interactions to segment the referred object throughout the video [23, 22, 68, 69, 70]. LRE [13] performs referred object localization and regional segmentation. URVOS [36] combines vision-language and temporal features for iterative segmentation. MMTBVS [71] uses motion clues to enhance multimodal representations. MLRL [72] extracts video-level, frame-level, and object-level features for comprehensive representations. Clawcranenet [73] and CMPC [74] mimic human behavior to progressively execute discriminative vision-language fusion. Similarly, HINet [75], YOFO [21], and LBDT [76] leverage multi-layer features for hierarchical cross-modal fusion.
Among these, transformer-based approaches have gained significant attention due to their end-to-end architecture and impressive performance. MTTR [24] introduces spatio-temporal transformer [77] for visual encoding and conditional kernels [25] for dynamic object segmentation. ReferFormer [23] employs language as queries to segment referred objects and includes an auxiliary detection loss to facilitate object localization learning. MANet [78] compresses video data and focuses on keyframe segmentation to speed up R-VOS. UNINEXT [79] introduces a unified model capable of training with data from various tasks. R2VOS [80] enhances multimodal alignment through text reconstruction. OnlineRefer [81] proposes an online model with explicit query propagation. HTML [82] performs multimodal interactions across various temporal scales. SgMg [22] proposes a segment-and-optimize paradigm to solve the feature drift issue. This work proposes an end-to-end paradigm that seamlessly incorporates temporal consistency modeling to achieve consistent and accurate R-VOS.
@@ III-C1 Multimodal Transformer
To generate object-related representations, cross-attention is employed to enhance each layer of the visual feature pyramid $\mathcal{F}^{v}$ by incorporating linguistic information $\mathcal{F}^{w}$,
@@
where $\odot$ and $W$ denote Hadamard product and learnable projection matrix. ${\rm Attn}(Q,K,V)$ represents the standard attention block [16] with query, key, and value. The multimodal transformer integrates information from the vision-language feature pyramid $\mathcal{F}^{vl}$ to effectively handle objects of varying sizes and scales. In this work, we adopt the memory-efficient Deformable Transformer [87] to encode vision-language features $\mathcal{F}^{vl}$ and predict object embeddings $\mathcal{E}$, similar to [23, 88, 89]. The multimodal transformer encoder produces a feature pyramid of the same size as the input. The transformer decoder uses a set of language-guided object queries $\mathcal{Q}$, generated by combining sentence features $\mathcal{F}^{s}$ and learnable embeddings, and cross-attends them with vision-language features $\mathcal{F}^{vl}$ to predict frame-wise object embeddings $\mathcal{E}$
@@ III-F Mask Consistency Score
Temporal consistency in video segmentation is crucial as it ensures that object masks remain accurate and consistent throughout the video, regardless of changes in appearance or context. Inconsistent segmentation can cause flickering and loss of tracked targets. However, existing R-VOS models primarily use frame-wise accuracy metrics for evaluation. To the best of our knowledge, there are currently no dedicated metrics designed to evaluate the temporal consistency of R-VOS models.
To fill this gap, we propose a new metric, Mask Consistency Score (MCS), to evaluate the temporal consistency of video segmentation at specified thresholds. The segmentation of a video is considered consistent if the accuracy (Jaccard index) for all frames exceeds the threshold $\tau$. Otherwise, the segmentation is deemed inconsistent due to the loss of tracking targets in one or more frames. The MCS metric is defined as the ratio of videos with consistent object segmentation. Let $\mathcal{J}_{x,y}$ denotes the Jaccard index of the segmentation for frame $y$ in video $x$, the MCS at threshold $\tau$ is formulated as
@@ III-G1 Loss for Selective Referring Segmentation
Selective segmentation predicts object masks $\mathcal{M}$, bounding boxes $\mathcal{B}$, and scores $\mathcal{S}$ for each video. Given a set of predictions $y=\{\forall y_{i},i\in[1,...,N]\}$ obtained from $N$=5 queries, where $y_{i}={\{\mathcal{M}}_{i,j},\mathcal{B}_{i,j},\mathcal{S}_{i,j}\}_{j=1}^{T}$, we employ the Hungarian algorithm [92] to find the optimal assignment with the highest similarity to the ground truth for training. The selective referring segmentation loss can be represented as
@@ III-G2 Loss for Inter-frame Collaboration
TABLE I: Quantitative comparison on Ref-YouTube-VOS and Ref-DAVIS17. : pre-trained with RefCOCO/g/+ images. S-HTR: integrated our hybrid memory with the recent work SgMg [22] for comparison. Our models achieve top-ranked performance in terms of accuracy ($\mathcal{J}$$&\mathcal{F}$) and temporally consistency (MCS@$\tau$)
@@ IV-A1 Datasets
We conduct experiments on four popular and challenging video datasets: Ref-YouTube-VOS [36], Ref-DAVIS17 [13], A2D-Sentences [14], JHMDB-Sentences [37]. Ref-YouTube-VOS is a large-scale dataset that consists of 3,471 training videos with 12,913 expressions and 507 validation videos with 2,096 expressions. Ref-DAVIS17 comprises 90 videos with 1,544 expressions, captured at 24 FPS. A2D-Sentences is an actor and action segmentation dataset with 3,782 video sequences, 6,656 action descriptions, and 1 to 5 frame annotations per sequence. JHMDB-Sentences includes 928 short video sequences and 928 descriptions covering 21 different action classes.
@@ IV-B Implementation Details
Following [22, 24, 23], we train our models on the Ref-YouTube-VOS train split and evaluate them on both Ref-YouTube-VOS and Ref-DAVIS17. We also provide results of our models that are pre-trained on static image datasets Ref-COCO/g/+ [100, 101] and then fine-tuned on Ref-YouTube-VOS. For A2D-Sentences and JHMDB-Sentences, we fine-tune the pre-trained models on the A2D-Sentences train split and evaluate them on both A2D-Sentences and JHMDB-Sentences.
We maintain the same optimization strategies as [23, 22] for a fair comparison. Our models are optimized using the AdamW [102] optimizer with a weight decay of 5×$\times 10^{-4}$. The initial learning rates are set to 2.5×$\times 10^{-5}$ for the visual encoder and inter-frame collaboration module, 5×$\times 10^{-6}$ for the textual encoder, and $\times 10^{-5}$ for the remaining components. The batch size is 4/2 for pre-training/main training and distributed on two 24GB RTX 3090 GPUs. During pre-training, each sample consists of one static image, we train our models for 12 epochs with a learning rate decay by a factor of 10 at the 8th and 10th epochs. During main training, each sample consists of 5 randomly sampled frames, we freeze the text encoder and train the models for 6/10 epochs with a learning rate decay by a factor of 10 at the 3/6th and 5/8th epochs, depending on whether pre-training was performed. The loss function coefficients, $\lambda_{dice}$, $\lambda_{L1}$, $\lambda_{focal}$, $\lambda_{GIoU}$, $\lambda_{ce}$ are set to 5, 5, 2, 2, 1, respectively. During inference, we follow [23, 22] to perform video-level segmentation ($N$=frame number) on Ref-YouTube-VOS and per-clip segmentation ($N$=36) on Ref-DAVIS17. We selectively segment the top 25% of ranked frames to construct memory for each video (see Table VI for the ablation study).
@@
Figure 4: Qualitative results. (a) Our HTR predicts temporally consistent results compared to ReferFormer [23] and SgMg [22]. (b) HTR can handle various challenging scenarios, including fast motion, objects with similar appearance, occlusion, and small objects. (c) HTR fails to improve mask quality without good reference masks.
@@ IV-C1 Results on Ref-YouTube-VOS
Table I shows that our HTR (with Swin-L) achieves the top-ranked performance of 67.1 $\mathcal{J}\&\mathcal{F}$ on Ref-YouTube-VOS, significantly outperforming the baseline ReferFormer [23] by 4.7% points. Without pre-training on Ref-COCO/g/+, HTR with VideoSwin-T achieves 59.8 $\mathcal{J}\&\mathcal{F}$ at 58 FPS, still surpassing ReferFormer while running about 1.2$\times$ faster. Despite its selective segmentation stage being built upon ReferFormer, HTR exceeds all recent competitors, including OnlineRefer [81], HTML [82], and SgMg [22], particularly in terms of temporal consistency. This demonstrates the effectiveness of our approach. Additionally, we integrate our hybrid memory with the recent work SgMg [22] to create the end-to-end model S-HTR for comparison. This integration further improves performance by a large margin. Without pre-training on Ref-COCO/g/+, S-HTR with VideoSwin-T achieves 61.5 $\mathcal{J}\&\mathcal{F}$, surpassing SgMg by 2.6% points. These results further demonstrate our architecture is effective and generic.

### v2

# Temporally Consistent Referring Video Object Segmentation with Hybrid Memory

@@ Abstract
Referring Video Object Segmentation (R-VOS) methods face challenges in maintaining consistent object segmentation due to temporal context variability and the presence of other visually similar objects. We propose an end-to-end R-VOS paradigm that explicitly models temporal instance consistency alongside the referring segmentation. Specifically, we introduce a novel hybrid memory that facilitates inter-frame collaboration for robust spatio-temporal matching and propagation. Features of frames with automatically generated high-quality reference masks are propagated to segment the remaining frames based on multi-granularity association to achieve temporally consistent R-VOS. Furthermore, we propose a new Mask Consistency Score (MCS) metric to evaluate the temporal consistency of video segmentation. Extensive experiments demonstrate that our approach enhances temporal consistency by a significant margin, leading to top-ranked performance on popular R-VOS benchmarks, i.e., Ref-YouTube-VOS (67.1%) and Ref-DAVIS17 (65.6%). The code is available at https://github.com/bo-miao/HTR.
@@ I Introduction
Early R-VOS methods [13, 14] primarily employ convolution networks [15] for visual encoding and language grounding. However, their inherent limitations in capturing long-range dependencies and handling free-form features result in suboptimal performance. With the advancement in attention mechanism [16, 17, 18, 19], various attention-based R-VOS approaches have emerged, leading to significant improvements. Techniques such as YOFO [20] and SgMg [21] employ cross-attention to model multimodal representations, enabling flexible and long-range vision-language correspondence. Furthermore, ReferFormer [22] and MTTR [23] introduce end-to-end transformers with object queries to dynamically separate target objects from semantic features.
These transformer-based techniques employ a set of language-guided object queries to attend to multimodal features of different frames, eventually predicting conditional kernels [24] for each frame. The frame-specific conditional kernels are then used to generate object masks by convolving the corresponding multimodal features. Despite demonstrating favorable performance, these approaches struggle to maintain consistent segmentation in complex and challenging scenarios. As illustrated in the first row of Fig. 1(b), the target sheep is accurately segmented in the initial frame, but the presence of other similar-looking sheep and changes in the local and global context, e.g., the disappearance of the dog in the final frames, lead to the failure of the model to consistently segment the correct target.
To mitigate this inconsistency issue, some approaches utilize well-trained Semi-VOS models [25, 26] for post-processing, e.g., ReferFormer [22] and PMINet [27] use CFBI [25] as a post-processing step to refine the masks. Such a naïve strategy requires a complex pipeline with additional backbones [15, 28] and training data [29, 30, 31, 32, 33, 34], making R-VOS cumbersome and impractical. Moreover, this strategy does not yield satisfactory improvements in performance as we show in Table V. An ideal solution would be an end-to-end paradigm that seamlessly integrates temporal consistency into R-VOS.
@@ II-A Semi-supervised Video Object Segmentation
Semi-VOS aims to segment target objects throughout the entire video, given the ground-truth mask (of the first frame) [37, 38, 39, 40]. Existing techniques mainly fall into three paradigms: finetuning-based, tracking-based, and matching-based. Finetuning-based methods dynamically update networks during test time using the first [41, 42], selected [43], or synthetic [44] frames from each video to make models object-specific. Despite their promising results, the slow test-time adaptation makes them inefficient. Tracking-based approaches [25, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54] avoid online learning and perform frame-to-frame propagation using masks of neighboring and annotated frames. However, they often lack sufficient long-term context, leading to the loss of targets after occlusions.
To address the context limitations, matching-based methods [55, 56, 57, 58, 59, 60, 61, 62, 12] construct memory banks with features from past frames and use attention mechanisms for temporal propagation. STM [63] and STCN [26] build memory banks to store mask and visual features at fixed intervals. SST [64] uses transformer blocks for affinity computation. AOT [11] and its subsequent works [65] establish hierarchical propagation to process multiple objects in parallel. XMem [10] introduces complementary memories inspired by the Atkinson-Shiffrin model [66] to efficiently handle long videos. In this work, we explore temporally consistent R-VOS, a task more challenging than Semi-VOS due to multimodal reasoning and the lack of ground-truth reference mask during inference, and propose an end-to-end paradigm based on a hybrid memory. Unlike STM [63] and STCN [26], our hybrid memory employs both local and global contexts for robust propagation of imperfect reference masks. In contrast to CFBI [25] and PMNet [67], the global tokens in our memory are constructed in one-stage and perform explicit node-object matching.
@@ II-B Referring Video Object Segmentation
R-VOS methods conduct vision-and-language interactions to segment the referred object throughout the video [22, 21, 68]. LRE [13] performs referred object localization and regional segmentation. URVOS [35] combines vision-language and temporal features for iterative segmentation. UMVOS [69] adopts reinforcement learning to optimize multi-modality filter networks. MMTBVS [70] uses motion clues to enhance multimodal representations. MLRL [71] extracts video-level, frame-level, and object-level features for comprehensive representations. Clawcranenet [72] and CMPC [73] mimic human behavior to progressively execute discriminative vision-language fusion. Similarly, HINet [74], YOFO [20], and LBDT [75] leverage multi-layer features for hierarchical cross-modal fusion.
Among these, transformer-based approaches have gained significant attention for their end-to-end architecture and impressive performance. MTTR [23] introduces spatio-temporal transformer [76] for visual encoding and conditional kernels [24] for dynamic object segmentation. ReferFormer [22] employs language as queries to segment referred objects and includes a detection loss to facilitate object localization learning. MANet [77] compresses video data and focuses on keyframe segmentation to speed up R-VOS. UNINEXT [78] introduces a unified model capable of training with data from various tasks. R2VOS [79] enhances multimodal alignment through text reconstruction. OnlineRefer [80] proposes an online model with explicit query propagation. HTML [81] performs multimodal interactions across various temporal scales. SgMg [21] proposes a segment-and-optimize paradigm to solve the feature drift issue. DsHmp [82] decouples video-level expression understanding into static and motion perception to enhance temporal comprehension. This work proposes an end-to-end paradigm that seamlessly incorporates temporal consistency modeling to achieve consistent and accurate R-VOS.
@@ III-C1 Multimodal Transformer
To generate object-related multimodal representation $\mathcal{F}^{vl}_{l}$, cross-attention is employed to enhance each layer of the visual feature pyramid $\mathcal{F}^{v}$ by incorporating linguistic information $\mathcal{F}^{w}$,
@@
where $\odot$ is the Hadamard product and $W$ represents learnable projection matrices. ${\rm Attn}(Q,K,V)$ is the standard attention block [16] with query, key, and value.
The multimodal transformer integrates information from the vision-language feature pyramid $\mathcal{F}^{vl}$ to effectively handle objects of varying sizes and scales. In this work, we adopt the memory-efficient Deformable Transformer [87] to encode vision-language features $\mathcal{F}^{vl}$ and predict object embeddings $\mathcal{E}$, similar to [22, 88, 89]. The transformer encoder takes $\mathcal{F}^{vl}$ as input and outputs a feature pyramid of the same size. The transformer decoder then uses a set of language-guided object queries $\mathcal{Q}$, generated by combining sentence features $\mathcal{F}^{s}$ and learnable embeddings, and cross-attends them with encoded vision-language features $\mathcal{F}^{vl}$ to predict frame-wise object embeddings $\mathcal{E}$,
@@ III-F Mask Consistency Score
Temporal consistency in video segmentation is crucial to maintain accurate and consistent object masks throughout the video, despite changes in appearance or context. Inconsistent segmentation can cause flickering and loss of tracked targets. However, existing R-VOS models primarily use frame-wise accuracy metrics for evaluation. To the best of our knowledge, there are currently no dedicated metrics designed to evaluate the temporal consistency of R-VOS models.
To fill this gap, we propose a new metric, Mask Consistency Score (MCS), to evaluate the temporal consistency of video segmentation at specified thresholds. The segmentation of a video is considered consistent if the accuracy (Jaccard index) for all frames exceeds the threshold $\tau$. Otherwise, the segmentation is deemed inconsistent due to the loss of tracking targets in one or more frames. The MCS metric is defined as the ratio of videos with consistent object segmentation. Different from mTC [92] for semantic segmentation, our metric does not require optical flow and imposes a strong penalty for the loss of targets (temporal inconsistency). Let $\mathcal{J}_{x,y}$ denotes the Jaccard index of the segmentation for frame $y$ in video $x$, the MCS at threshold $\tau$ is formulated as
@@ III-G1 Loss for Selective Referring Segmentation
Selective segmentation predicts object masks $\mathcal{M}$, bounding boxes $\mathcal{B}$, and scores $\mathcal{S}$ for each video. Given a set of predictions $y=\{\forall y_{i},i\in[1,...,N]\}$ obtained from $N$=5 object queries, where $y_{i}={\{\mathcal{M}}_{i,j},\mathcal{B}_{i,j},\mathcal{S}_{i,j}\}_{j=1}^{T}$, we employ the Hungarian algorithm [93] to find the optimal assignment with the highest similarity to the ground truth for training. The selective referring segmentation loss can be represented as
@@ III-G2 Loss for Inter-frame Collaboration
TABLE I: Quantitative comparison on Ref-YouTube-VOS and Ref-DAVIS17. : pre-trained with RefCOCO/g/+ images. : uses 20 queries and trains for 30 epochs. S-HTR: integrates our hybrid memory with the recent work SgMg [21] for comparison. Our models achieve top-ranked performance in terms of accuracy ($\mathcal{J}$$&\mathcal{F}$) and temporally consistency (MCS@$\tau$)
@@ IV-A1 Datasets
We conduct experiments on four popular and challenging video datasets: Ref-YouTube-VOS [35], Ref-DAVIS17 [13], A2D-Sentences [14], JHMDB-Sentences [36], and MeViS [100]. Ref-YouTube-VOS is a large-scale dataset that consists of 3,471 training videos with 12,913 expressions and 507 validation videos with 2,096 expressions. Ref-DAVIS17 comprises 90 videos with 1,544 expressions, captured at 24 FPS. A2D-Sentences is an actor and action segmentation dataset with 3,782 video sequences, 6,656 action descriptions, and 1 to 5 frame annotations per sequence. JHMDB-Sentences includes 928 short video sequences and 928 descriptions covering 21 different action classes.
@@ IV-B Implementation Details
Following [21, 23, 22], we train our models on the Ref-YouTube-VOS train split and evaluate them on both Ref-YouTube-VOS and Ref-DAVIS17. We also present results for our models pre-trained on static image datasets Ref-COCO/g/+ [102, 103] and fine-tuned on Ref-YouTube-VOS. For A2D-Sentences and JHMDB-Sentences, we fine-tune the pre-trained models on the A2D-Sentences train split and evaluate them on both A2D-Sentences and JHMDB-Sentences.
We maintain the same optimization strategies as [22, 21] for fair comparison. Our models are optimized with AdamW [104] and a weight decay of 5×$\times 10^{-4}$. The initial learning rates are set to 2.5×$\times 10^{-5}$ for the visual encoder and inter-frame collaboration module, 5×$\times 10^{-6}$ for the textual encoder, and $\times 10^{-5}$ for the rest. The batch size is 4/2 for pre-training/main training and distributed on two 24GB RTX 3090 GPUs. During pre-training, each sample consists of one static image, we train our models for 12 epochs with a learning rate decay by a factor of 10 at the 8th and 10th epochs. During main training, each sample consists of 5 randomly sampled frames, we freeze the text encoder and train the models for 6/10 epochs with a learning rate decay by a factor of 10 at the 3/6th and 5/8th epochs, depending on whether pre-training was performed. The loss function coefficients, $\lambda_{dice}$, $\lambda_{L1}$, $\lambda_{focal}$, $\lambda_{GIoU}$, $\lambda_{ce}$ are set to 5, 5, 2, 2, 1, respectively. During inference, we follow [22, 21] to perform video-level segmentation ($N$=frame number) on Ref-YouTube-VOS and per-clip segmentation ($N$=36) on Ref-DAVIS17. We selectively segment the top 25% of ranked frames to construct memory for each video (see Table IX for the ablation study).
@@
Figure 4: Qualitative results on Ref-YouTube-VOS and MeViS. (a) Our HTR predicts temporally consistent results compared to ReferFormer [22] and SgMg [21]. (b) HTR can handle appearance and motion expressions in various challenging scenarios, including fast motion, objects with similar appearance, occlusion, and small objects. (c) HTR fails to improve mask quality without good reference masks.
@@ IV-C1 Results on Ref-YouTube-VOS
Table I shows that our HTR (with Swin-L) achieves the top-ranked performance of 67.1 $\mathcal{J}\&\mathcal{F}$ on Ref-YouTube-VOS, significantly outperforming the baseline ReferFormer [22] by 4.7% points. Without pre-training on Ref-COCO/g/+, HTR with VideoSwin-T achieves 59.8 $\mathcal{J}\&\mathcal{F}$ at 58 FPS, still surpassing ReferFormer while running about 1.2$\times$ faster. Despite its selective segmentation stage being built upon ReferFormer, HTR exceeds all recent competitors, including OnlineRefer [80], HTML [81], SgMg [21], and SOC [98], particularly in terms of temporal consistency. This demonstrates the effectiveness of our approach. Additionally, we integrate our hybrid memory with the recent work SgMg [21] to create the end-to-end model S-HTR for comparison. This integration further improves performance by a large margin. Without pre-training on Ref-COCO/g/+, S-HTR with VideoSwin-T achieves 61.5 $\mathcal{J}\&\mathcal{F}$, surpassing SgMg by 2.6% points. These results further demonstrate our architecture is effective and generic.